How many ounces of 25% alcohol solution must be mixed with 18 ounces of 30% solution to make 27% alcohol?

Answer by mananth(16946)

--------percent ---------------- quantity
Alcohol I 30 ---------------- 18 Oz
alcohol II 25 ---------------- x Oz
Mixture 27 ---------------- 18 + x Oz

30*18+25x=27(18+x)
540+25x =486+27 x
25x-27x =486+540
-2x=-54
/ -2
x = 27
27 Oz of 25% alcohol
m.ananth@hotmail.ca